# Context-Free Recognition with Transformers

### v1

@@ 4.1 A path system framework for unambiguous CFL recognition
${\mathcal{E}}={{\{({{{z}}},{{{x}}})\mid{{{z}}}\notin{{{\mathcal{X}}}},{\mathcal{R}({{{x}}},{{{y}}},{{{z}}})}={\mathrm{T}}\text{ or }{\mathcal{R}({{{y}}},{{{x}}},{{{z}}})}={\mathrm{T}}\text{ for some }{{{y}}}\in{{{\mathcal{X}}}}\}}}$ (1)
@@
1.def $\textsc{Compute closure}({{\bm{w}}},{{\mathcal{G}}})$:
2.initialize ${\mathcal{V}}\leftarrow{{\{[{{\mathrm{A}}},{{i}},{{j}}]\}}}$
3.initialize ${\mathcal{T}}\leftarrow{{\{[{{\mathrm{A}}},{{i}},{{i}}]\mid{{{{\mathrm{A}}}\rightarrow{\texttt{w}}_{{i}}}}\in{{\mathcal{P}}}\}}}$
@@
5.for _ in range $\log({{n}})$ :
6.  ${{{\mathcal{D}}}}\leftarrow{\textsc{DG}}({{{\mathcal{X}}}})$
7.  ${{{\mathcal{X}}}}\leftarrow\textsc{Reach}({{{\mathcal{D}}}})$
8.return $\mathcal{X}$}}}
Algorithm 3 Algorithm for computing ${\mathcal{C}}({{\bm{w}}})$
@@ 5 Experiments
We conduct experiments to elicit the impact of looping when recognizing formal languages, and provide more details on our experimental setup in App.˜C. We train transformer classifiers on CFLs of varying degrees of complexity:
  • Boolean formula value problem (BFVP): The set of variable-free Boolean formulas that evaluate to $\mathrm{T}$}. We have proven (Cor.˜4.1) that $\log$-depth and no padding suffice to recognize this language on transformers. We consider formulas in both the standard infix notation (e.g., $1\lor 0$ is in infix notation) and postfix notation (e.g., 11 0 $\lor$ is in postfix notation). Parallel algorithms for BFVP typically rely on postfix notation (buss-bfvp; buss-optimalparallel).
@@
These languages vary in complexity, allowing us to test transformers’ ability to learn CFL recognition constructions for languages of different difficulties. In particular, while Palindrome and $\mathrm{D}\!\left(k\right)$}} languages can in principle be recognized by constant-depth transformers, BFVP requires growing depth (i.e., $\log$-depth), assuming ${\textsc{TC}\textsuperscript{0}}\neq{\textsc{NC}\textsuperscript{1}}$. This suggests that the performance of log-depth vs. constant-depth transformers on BFVP is a good measure of whether transformers can utilize the extra expressivity of log-depth when it is required.
Language Fixed-depth $\log({{n}})$ looping Fixed-depth $\log({{n}})$ looping
BFVP $0.97_{\pm_{0}.01}$ $0.98_{\pm 0.00}$ $0.88_{\pm 0.01}$ $0.91_{\pm 0.01}$
Palindrome $0.94_{\pm 0.01}$ $0.93_{\pm 0.01}$ $0.79_{\pm 0.03}$ $0.72_{\pm 0.03}$
Results.
We first highlight that for both variants of BFVP, looping leads to slight improvements in in-distribution (1-3%) and generalization (3-4%) accuracy. We treat BFVP as a testbed of our theory as it is a language known to require $\log$-depth. As we see improvements in performance with looping, our results align with the theory. However, the results are mixed for other languages. For Palindrome and $\mathrm{D}\!\left(1\right)$}}, we notice looping does not improve accuracy. This is supported by the fact that these languages already have fixed-size solutions (hao-etal-2022-formal; weiss2021thinkingliketransformers), and therefore extending the model with dynamically-scaling layers may hinder the performance. However, for $\mathrm{D}\!\left(2\right)$}} and Marked Palindrome, we remark that looping seems to improve generalization even though these languages also have constant-depth transformer constructions. This is supported by the fact that these languages are inexpressible in C-RASP (yang2024countingliketransformerscompiling; huang2025a), a language class that matches closely the set of languages transformers should be able to generalize on.
@@ 6 Discussion and conclusion
We show that transformers with $\log$-depth can recognize general CFLs if they can use padding tokens (merrill2025exactexpressivepowertransformers). In addition, we characterize unambiguity and linearity as CFL features that can reduce the amount of padding needed by transformers for recognition. These results reveal one way that transformers with limited depth can recognize CFLs and predict ambiguity in language could be a hurdle for transformers to process, as suggested in previous empirical work (khalighinejad-etal-2023-approximating; liu2023wereafraidlanguagemodels). Although it is not possible to improve our $\log$-depth recognition algorithm to fixed depth unless ${\textsc{TC}\textsuperscript{0}}={\textsc{NC}\textsuperscript{1}}$, our padding bounds are not known to be tight. Therefore, future work could find more padding-efficient transformer constructions for recognizing general CFLs, or subclasses thereof. Finally, as expressivity results cannot fully predict the empirical abilities of transformers, recent learnability results (hahn-rofin-2024-sensitive) are painting a more complete picture of the abilities and limitations of transformers. We therefore encourage future work to investigate theoretically the conditions under which a transformer can learn to process syntax on out-of-distribution inputs.
@@ Acknowledgments
Anej Svete is supported by an ETH AI Center Doctoral Fellowship. William Merrill was supported by a Two Sigma PhD fellowship, an NSF Graduate Research Fellowship, and the Allen Institute for Artifical Intelligence.
@@ A.1.1 Fixed-size transformers
${{\textsf{T}}}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\mathrm{\mathcal{L}}}^{({L})}\circ\cdots\circ{\mathrm{\mathcal{L}}}^{(1)}\circ{\texttt{embed}}$ (2)
@@
${\mathrm{\mathcal{L}}}^{({\ell})}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\mathbf{F}}^{({\ell})}\circ{{{\bm{f}}}_{\texttt{att}}}^{({\ell})}\circ{\texttt{LN}}^{({\ell})}$ (3)
We recall layer-normalization maps a vector ${{{\bm{x}}}}\in{{\mathbb{R}}}^{{n}}$ of some dimension n{{n}} to $\frac{{{{\bm{x}}}}^{\prime}}{{{\left\lVert{{{\bm{x}}}}^{\prime}\right\rVert}}}$ where ${{{\bm{x}}}}^{\prime}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{{{\bm{x}}}}-\frac{\sum_{{{{x}}}_{{i}}\in{{{\bm{x}}}}}{{{x}}}_{{i}}}{{{n}}}$. We assume multi-pre-norm (merrill2024expressivepowertransformerschain). In standard pre-norm, we apply a layer-normalization to the entire hidden state of each symbol. In multi-pre-norm, we allow each sublayer to take $k$ different projections of its input apply layer-norm to each and concatenate. Crucially, multi-pre-norm allows us to partition the hidden state and normalize disjoint subsets thereof, which we will rely on in our proofs.
@@
${{{\bm{f}}}_{\texttt{att}}}^{({\ell})}(({{{x}}}_{1},\cdots,{{{x}}}_{{{T}}}))$ $\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}({{{y}}}_{1},\cdots,{{{y}}}_{{{T}}})$ (4a)
${{{y}}}_{{i}}$ $\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{{{x}}}_{{i}}+\sum_{{{i}}^{\prime}\in{{m}}\left({{i}}\right)}{{{s}}}_{{{i}}^{\prime}}{{{\bm{v}}}}_{{{i}}^{\prime}}^{({\ell})}$ (4b)
@@
${{m}}\left({{i}}\right)$ is a set that defines the masking used by the transformer. For instance, ${{m}}\left({{i}}\right)={{\{{{i}}^{\prime}\mid{{i}}^{\prime}<{{i}}\}}}$ refers to strict causal masking and ${{m}}\left({{i}}\right)={{\left[|{{\bm{w}}}|\right]}}$ refers to no masking. score is a scoring function that maps two vectors of the same size to a scalar. Typically, the dot-product score is used with ${{{\texttt{score}}}}({{{x}}}_{1},{{{x}}}_{2})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\langle{{{x}}}_{1},{{{x}}}_{2}\rangle$.
@@
for ${{d}}\in{{\left[{{D}}\right]}}$, where ${{{\bm{x}}}}\in{{\mathbb{R}}}^{{D}}$ and $m\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}|\operatorname*{{{argmax}}}\left({{{\bm{x}}}}\right)|$ is the cardinality of the argmax set.
@@
${{\mathbb{L}}}\left({{\textsf{T}}}\right)\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{{\{{{\bm{w}}}\in{{{{\Sigma}}^{*}}}\mid{{{\bm{\theta}}}}^{\top}{{{\bm{x}}}}_{{{\textsc{eos}}}}^{{L}}>0\}}}.$ (6)
@@ A.1.2 Layer-norm hash
Given a scalar ${{{z}}}\in\mathbb{R}$, its layer-norm hash is $\phi({{{z}}})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\langle{{{z}}},1,-{{{z}}},-1\rangle/\sqrt{{{{z}}}^{2}+1}$.
@@ B.1 General CFL recognition on transformers
Items of the form ${{\mathrm{X}}},{{i}},{{j}}]$ are a base case item if i=j{{i}}={{j}}. A feedforward network can for each padding token associated with some ${{\mathrm{X}}},{{i}},{{j}}]$ check that i=j{{i}}={{j}} by adding ij{{i}}-{{j}} to the residual stream. With an attention layer, we can then retrieve and add to the residual stream the encoding of the symbol ${\texttt{w}}_{{i}}$ for a given base case item ${{\mathrm{X}}},{{i}},{{i}}]$ as follows. A symbol representation at position i{{i}} can add to its residual stream $\phi({{i}})$ by uniformly attending with a causally-masked attention layer to all symbol representations in the strict left context (merrill2024expressivepowertransformerschain). A padding token associated with ${{\mathrm{X}}},{{i}},{{i}}]$ also stores $\phi({{i}})$. Therefore, via an equality-check via dot product, padding tokens can attend to relevant symbol representations by setting as value the one-hot encoding of the symbol $\llbracket{\texttt{w}}_{{i}}\rrbracket$}}. Finally, a feedforward network can add to the residual stream 1 if ${{\mathrm{X}}}\rightarrow{\texttt{w}}_{{i}}$}} is a valid rule and otherwise 0: A mapping between two finite sets ${{\mathcal{N}}}\times{{\Sigma}}\to{{\{0,1\}}}$ can be computed by a feedforward network.
@@ B.2 Unambiguous CFL recognition on transformers
  • If $\alpha$ and $\beta$ are a formulas in postfix notation then $\alpha\beta\land$ and $\alpha\beta\lor$ are formulas in postfix notation.
We first detail how a transformer can determine whether an input formula is well-formed and how it can add to the residual streams of tokens representing an operator $\lor$ (or $\land$, $\neg$) the encodings to the arguments of that operator. To this extent, we will write a C-RASP program (yang2024countingliketransformerscompiling) to compute whether an input formula is well-formed and what are the arguments of each operator in the formula. C-RASP defines a syntax for writing programs, and a C-RASP program defines a formal language. It is known that $\text{{C-RASP}}\subseteq{{\textsc{AHAT}}}^{0}_{0}$ (yang2024countingliketransformerscompiling).
@@
A padding token for some node can check whether its associated item is of the form ${{\mathrm{A}}},{{i}},{{i}}]$ via a feedforward network that checks that the indices are the same. For all such padding tokens, another feedforward network adds $1$ to the residual stream if and only if ${{{{\mathrm{A}}}\rightarrow{\texttt{w}}_{{i}}}}\in{{\mathcal{P}}}$ to signal the realizability of that item (and otherwise adds 0). We can perform this procedure exactly as in the base case of §˜B.1.
Creating the dependency graph:
@@
Padding tokens for edges store items of the form ${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$. There are finitely many ${{\mathrm{C}}},{{k}}+1,{{j}}]$ such that ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$ (proportionally many in $|{{\mathcal{N}}}|$), which can be added to the residual stream via a feedforward network. According to Eq.˜1, we set an edge between nodes associated with ${{\mathrm{A}}},{{i}},{{j}}]$ and ${{\mathrm{B}}},{{i}},{{k}}]$ if and only if there is an item ${{\mathrm{C}}},{{k}}+1,{{j}}]$ such that ${{\mathrm{C}}},{{k}}+1,{{j}}]$ is realizable (i.e, the corresponding padding token stores 1 in its residual stream) and ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$. The padding token for the edge associated with ${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$ can check whether any of the items of the form ${{\mathrm{C}}},{{k}}+1,{{j}}]$ are realizable and satisfies ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$ via an equality-check with an attention layer (to check the realizability of the items) and a feedforward-network (to check whether ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$). If such an item exists, the padding token associated with ${{\mathrm{A}}},{{i}},{{j}}]$ and ${{\mathrm{B}}},{{i}},{{k}}]$ signals that there is an edge between them in the dependency graph.
Binarization:
@@
To efficiently perform reachability queries on a dependency graph, we require it to be binary (rytter-2pda). To this extent, we define the graph transform ${{\mathcal{T}}}\colon{\mathcal{G}}\to{\mathcal{G}}^{\prime}$ which binarizes a given directed graph $\mathcal{G}$} by adding more edges and nodes. Denoting ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ and ${\mathcal{G}}^{\prime}=({\mathcal{V}}^{\prime},{\mathcal{E}}^{\prime})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{{\mathcal{T}}}({\mathcal{G}})$, our graph transform satisfies ${\mathcal{V}}\subset{\mathcal{V}}^{\prime}$ as it simply adds more nodes to the original graph. We now describe $\mathcal{T}$}}.
@@
Given a node with k{{k}} out-neighbors, $\mathcal{T}$}} adds k{{k}}-2 extra nodes in $\mathcal{G}$$^{\prime}$ to create a right-branching binary tree. We denote by $r$ some root node and $v_{1},v_{2},\ldots v_{{k}}$ the out-neighbors of $v$ from $\mathcal{G}$}. $\mathcal{T}$}} introduces k{{k}}-2 extra nodes, $h_{1},h_{2},\ldots,h_{{{k}}-2}$, that are used as follows. The root node $r$ in $\mathcal{G}$$^{\prime}$ has edges to $v_{1}$ and $h_{1}$. For 0<i<k20<{{i}}<{{k}}-2, we instantiate the edges $(h_{{{n}}-2},v_{{{n}}-1})$ and $(h_{{{n}}-2},v_{{{n}}})$. For i=n2{{i}}={{n}}-2, we instantiate the edges $(h_{{{n}}-2},v_{{{n}}-1})$ and $(h_{{{n}}-2},v_{{{n}}})$. The resulting tree is a binary right-branching tree. Crucially, we get the following fact.
Fact B.1.
If there is a path between two nodes $v_{1},v_{2}$ in $\mathcal{G}$}, there is a path between the corresponding nodes $v_{1},v_{2}$ in ${{\mathcal{T}}}({\mathcal{G}})$.
Our construction trivially preserves reachability query results. We now show how to perform $\mathcal{T}$}} on transformers.

### v2

@@ 4.1 A Path System Framework for Unambiguous CFL Recognition
${\mathcal{E}}$ $=\{({{{z}}},{{{x}}})\mid{{{z}}}\notin{{{\mathcal{X}}}},{\mathcal{R}({{{x}}},{{{y}}},{{{z}}})}={\mathrm{T}}$ (1)
$\text{ or }{\mathcal{R}({{{y}}},{{{x}}},{{{z}}})}={\mathrm{T}}\text{ for some }{{{y}}}\in{{{\mathcal{X}}}}\}$ (2)
@@
$\textsc{Compute closure}({{\bm{w}}},{{\mathcal{G}}})$${\mathcal{V}}\leftarrow{{\{[{{\mathrm{A}}},{{i}},{{j}}]\}}}initialize$${\mathcal{T}}\leftarrow{{\{[{{\mathrm{A}}},{{i}},{{i}}]\mid{{{{\mathrm{A}}}\rightarrow{\texttt{w}}_{{i}}}}\in{{\mathcal{P}}}\}}}$${{{\mathcal{X}}}}\leftarrow{\mathcal{T}}$_ in range $\log({{n}})$${{{\mathcal{D}}}}\leftarrow{\textsc{DG}}({{{\mathcal{X}}}})$${{{\mathcal{X}}}}\leftarrow\textsc{Reach}({{{\mathcal{D}}}})$$\mathcal{X}$}}}
@@
[…]
@@
Language Fixed-depth $\log({{n}})$ looping Fixed-depth $\log({{n}})$ looping
BFVP $0.97_{\pm 0.01}$ $0.98_{\pm 0.00}$ $0.88_{\pm 0.01}$ $0.91_{\pm 0.01}$
Palindrome $0.94_{\pm 0.01}$ $0.93_{\pm 0.01}$ $0.79_{\pm 0.03}$ $0.72_{\pm 0.03}$
@@ 5 Experiments
We conduct experiments to elicit the impact of looping when recognizing CFLs, and provide more details on our experimental setup in App.˜C. We train transformer classifiers on CFLs of varying degrees of complexity, none of which require extra padding for recognition:
  • Boolean formula value problem (BFVP): The set of variable-free Boolean formulas that evaluate to $\mathrm{T}$}. We have proven (Cor.˜4.1) that $\log$-depth and no padding suffice to recognize this language with transformers. We consider formulas in both the standard infix notation (e.g., $1\lor 0$ is in infix notation) and postfix notation (e.g., 11 0 $\lor$ is in postfix notation). Parallel algorithms for BFVP typically rely on postfix notation (buss-bfvp; buss-optimalparallel).
@@
These languages vary in complexity, allowing us to test transformers’ ability to learn CFL recognition constructions for languages of different difficulties. In particular, while Palindrome and $\mathrm{D}\!\left(k\right)$}} languages can in principle be recognized by constant-depth transformers, BFVP requires growing depth (i.e., $\log$-depth), assuming ${\textsc{TC}\textsuperscript{0}}\neq{\textsc{NC}\textsuperscript{1}}$. This suggests that the performance of log-depth vs. constant-depth transformers on BFVP is a good measure of whether transformers can utilize the extra expressivity of log-depth when it is required. Our results are presented in Tab.˜2.
Results. We first highlight that for both variants of BFVP, looping leads to slight improvements in in-distribution (1-3%) and generalization (3-4%) accuracy. We treat BFVP as a testbed of our theory as it is a language known to require $\log$-depth. As we see improvements in performance with looping, our results align with the theory. However, the results are mixed for other languages. For Palindrome and $\mathrm{D}\!\left(1\right)$}}, we notice looping does not improve accuracy. This is supported by the fact that these languages already have fixed-size solutions (hao-etal-2022-formal; weiss2021thinkingliketransformers), and therefore extending the model with dynamically-scaling layers may hinder the performance. However, for $\mathrm{D}\!\left(2\right)$}} and Marked Palindrome, we remark that looping seems to improve generalization even though these languages also have constant-depth transformer constructions. This is supported by the fact that these languages are inexpressible in C-RASP (yang2024countingliketransformerscompiling; huang2025a), a language class that matches closely the set of languages transformers should be able to generalize on.
@@ 6 Discussion and Conclusion
Transformers parse in parallel.
Our work provides a theoretical framework for understanding how transformers can internally process syntax: The parsing problem can be formulated as a parallel procedure implementable by looped- and padded-transformers (§˜3, §˜4). Interestingly, transformers in practice seem to implement some form of parallel parsing. schulz2025unravelingsyntaxlanguagemodels show that transformers parse by learning sub-grammars—grammars that generate substrings of the original grammar—in parallel.
Most significantly, allen-zhu2025physics show via probing that transformers simulate a dynamic program that manipulates items: They found that transformers encode the necessary items (of the form ${{\mathrm{X}}},{{i}},{{j}}]$) to parse a string, and that they implement memory reads across positions to combine the solutions of items. While they state that such an algorithm can be naively implemented in polynomial time, our constructions leverage transformers’ inherent parallelism to show it can be implemented exponentially faster (i.e., in logarithmic time). Akin to how transformers can implement state tracking via parallel simulation of automata (liu2023transformerslearnshortcutsautomata; li2025how), our work provides a rigorous construction for how looped transformers can implement parallel parsing which complements recent interpretability results.
On dynamically-scaling transformers.
Fixed-size transformers cannot robustly solve the general CFL recognition problem, neither theoretically (merrill-etal-2022-saturated) nor empirically (khalighinejad-etal-2023-approximating; butoi-languageclasses). There is a simple explanation for this phenomenon: general CFL recognition is an NC1-hard problem (venkateswaran-logcfl). While transformers used in practice are fixed-size, our theoretical analysis can guide novel model architectures with improved algorithmic capabilities over any input.
Learnability.
When aiming to predict the exact empirical capabilities of some parametrized model, expressivity results cannot paint the full picture. A function that is expressible may not be easily learnable (hahn-rofin-2024-sensitive). While our work provides a framework for understanding how transformers can express a context-free language, we encourage future work to further investigate which class of grammars are provably learnable (huang2025a), and if looping and padding can enhance learnability (fan2025loopedtransformerslengthgeneralization).
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Acknowledgments
We thank attendees of Formal Languages and Neural Networks (FLaNN) seminar for insightful discussions about this work. Anej Svete is supported by an ETH AI Center Doctoral Fellowship. William Merrill was supported by a Two Sigma PhD fellowship, an NSF Graduate Research Fellowship, and the Allen Institute for Artificial Intelligence.
@@ A.1.1 Fixed-size Transformers
${{\textsf{T}}}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{\mathrm{\mathcal{L}}}^{({L})}\circ\cdots\circ{\mathrm{\mathcal{L}}}^{(1)}\circ{\texttt{embed}}$ (3)
@@
${\mathrm{\mathcal{L}}}^{({\ell})}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{\mathbf{F}}^{({\ell})}\circ{{{\bm{f}}}_{\texttt{att}}}^{({\ell})}\circ{\texttt{LN}}^{({\ell})}$ (4)
We recall layer-normalization maps a vector ${{{\bm{x}}}}\in{{\mathbb{R}}}^{{n}}$ of some dimension n{{n}} to $\frac{{{{\bm{x}}}}^{\prime}}{{{\left\lVert{{{\bm{x}}}}^{\prime}\right\rVert}}}$ where ${{{\bm{x}}}}^{\prime}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{{{\bm{x}}}}-\frac{\sum_{{{{x}}}_{{i}}\in{{{\bm{x}}}}}{{{x}}}_{{i}}}{{{n}}}$. We assume multi-pre-norm (merrill2024expressivepowertransformerschain). In standard pre-norm, we apply a layer-normalization to the entire hidden state of each symbol. In multi-pre-norm, we allow each sublayer to take $k$ different projections of its input apply layer-norm to each and concatenate. Crucially, multi-pre-norm allows us to partition the hidden state and normalize disjoint subsets thereof, which we will rely on in our proofs.
@@
${{{\bm{f}}}_{\texttt{att}}}^{({\ell})}(({{{x}}}_{1},\cdots,{{{x}}}_{{{T}}}))$ $\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}({{{y}}}_{1},\cdots,{{{y}}}_{{{T}}})$ (5a)
${{{y}}}_{{i}}$ $\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{{{x}}}_{{i}}+\sum_{{{i}}^{\prime}\in{{m}}\left({{i}}\right)}{{{s}}}_{{{i}}^{\prime}}{{{\bm{v}}}}_{{{i}}^{\prime}}^{({\ell})}$ (5b)
@@
${{m}}\left({{i}}\right)$ is a set that defines the masking used by the transformer. For instance, ${{m}}\left({{i}}\right)={{\{{{i}}^{\prime}\mid{{i}}^{\prime}<{{i}}\}}}$ refers to strict causal masking and ${{m}}\left({{i}}\right)={{\left[|{{\bm{w}}}|\right]}}$ refers to no masking. score is a scoring function that maps two vectors of the same size to a scalar. Typically, the dot-product score is used with ${{{\texttt{score}}}}({{{x}}}_{1},{{{x}}}_{2})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\langle{{{x}}}_{1},{{{x}}}_{2}\rangle$.
@@
for ${{d}}\in{{\left[{{D}}\right]}}$, where ${{{\bm{x}}}}\in{{\mathbb{R}}}^{{D}}$ and $m\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}|\operatorname*{{{argmax}}}\left({{{\bm{x}}}}\right)|$ is the cardinality of the argmax set.
@@
${{\mathbb{L}}}\left({{\textsf{T}}}\right)\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{{\{{{\bm{w}}}\in{{{{\Sigma}}^{*}}}\mid{{{\bm{\theta}}}}^{\top}{{{\bm{x}}}}_{{{\textsc{eos}}}}^{{L}}>0\}}}.$ (7)
@@ A.1.2 Layer-norm hash
Given a scalar ${{{z}}}\in\mathbb{R}$, its layer-norm hash is $\phi({{{z}}})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\langle{{{z}}},1,-{{{z}}},-1\rangle/\sqrt{{{{z}}}^{2}+1}$.
@@ B.1 General CFL Recognition on Transformers
Items of the form ${{\mathrm{X}}},{{i}},{{j}}]$ are a base case item if i=j{{i}}={{j}}. A feedforward network can for each padding token associated with some ${{\mathrm{X}}},{{i}},{{j}}]$ check that i=j{{i}}={{j}} by adding ij{{i}}-{{j}} to the residual stream and checking ij=0{{i}}-{{j}}=0 using $\mathrm{ReLU}$}} activations [^6] This equality check only works because ij{{i}}-{{j}} is guaranteed to be an integer (yang2025transformercookbook).. With an attention layer, we can then retrieve and add to the residual stream the encoding of the symbol ${\texttt{w}}_{{i}}$ for a given base case item ${{\mathrm{X}}},{{i}},{{i}}]$ as follows. A symbol representation at position i{{i}} can add to its residual stream $\phi({{i}})$ by uniformly attending with a causally-masked attention layer to all symbol representations in the strict left context (merrill2024expressivepowertransformerschain). A padding token associated with ${{\mathrm{X}}},{{i}},{{i}}]$ also stores $\phi({{i}})$. Therefore, via an equality-check via dot product, padding tokens can attend to relevant symbol representations by setting as value the one-hot encoding of the symbol $\llbracket{\texttt{w}}_{{i}}\rrbracket$}}. Finally, a feedforward network can add to the residual stream 1 if ${{\mathrm{X}}}\rightarrow{\texttt{w}}_{{i}}$}} is a valid rule and otherwise 0: A mapping between two finite sets ${{\mathcal{N}}}\times{{\Sigma}}\to{{\{0,1\}}}$ can be computed by a feedforward network.
@@ B.2 Unambiguous CFL Recognition on Transformers
  • If $\alpha$ and $\beta$ are formulas in postfix notation then $\alpha\beta\land$ and $\alpha\beta\lor$ are formulas in postfix notation.
We first detail how a transformer can determine whether an input formula is well-formed and how it can add to the residual streams of tokens representing an operator $\lor$ (or $\land$, $\neg$) the encodings to the arguments of that operator. To this extent, we will write a C-RASP program (yang2024countingliketransformerscompiling) to compute whether an input formula is well-formed and what are the arguments of each operator in the formula. C-RASP defines a syntax for writing programs, and a C-RASP program defines a formal language. It is known that C-RASP is equivalent to temporal counting logic (yang2024countingliketransformerscompiling), which is a lower bound on the expressive power of fixed-depth average-hard attention transformers (barcelo2024logical). It is therefore established that $\text{{C-RASP}}\subseteq{{\textsc{AHAT}}}^{0}_{0}$.
@@
Initial items. A padding token for some node can check whether its associated item is of the form ${{\mathrm{A}}},{{i}},{{i}}]$ via a feedforward network that checks that the indices are the same. For all such padding tokens, another feedforward network adds $1$ to the residual stream if and only if ${{{{\mathrm{A}}}\rightarrow{\texttt{w}}_{{i}}}}\in{{\mathcal{P}}}$ to signal the realizability of that item (and otherwise adds 0). We can perform this procedure exactly as in the base case of §˜B.1.
@@
Creating the dependency graph. Padding tokens for edges store items of the form ${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$. There are finitely many ${{\mathrm{C}}},{{k}}+1,{{j}}]$ such that ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$ (proportionally many in $|{{\mathcal{N}}}|$), which can be added to the residual stream via a feedforward network. According to Eq.˜1, we set an edge between nodes associated with ${{\mathrm{A}}},{{i}},{{j}}]$ and ${{\mathrm{B}}},{{i}},{{k}}]$ if and only if there is an item ${{\mathrm{C}}},{{k}}+1,{{j}}]$ such that ${{\mathrm{C}}},{{k}}+1,{{j}}]$ is realizable (i.e, the corresponding padding token stores 1 in its residual stream) and ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$. The padding token for the edge associated with ${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$ can check whether any of the items of the form ${{\mathrm{C}}},{{k}}+1,{{j}}]$ are realizable and satisfies ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$ via an equality-check with an attention layer (to check the realizability of the items) and a feedforward-network (to check whether ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$). If such an item exists, the padding token associated with ${{\mathrm{A}}},{{i}},{{j}}]$ and ${{\mathrm{B}}},{{i}},{{k}}]$ signals that there is an edge between them in the dependency graph.
@@
Binarization. To efficiently perform reachability queries on a dependency graph, we require it to be binary (rytter-2pda). To this extent, we define the graph transform ${{\mathcal{T}}}\colon{\mathcal{G}}\to{\mathcal{G}}^{\prime}$ which binarizes a given directed graph $\mathcal{G}$} by adding more edges and nodes. Denoting ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ and ${\mathcal{G}}^{\prime}=({\mathcal{V}}^{\prime},{\mathcal{E}}^{\prime})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{{\mathcal{T}}}({\mathcal{G}})$, our graph transform satisfies ${\mathcal{V}}\subset{\mathcal{V}}^{\prime}$ as it simply adds more nodes to the original graph. We now describe $\mathcal{T}$}}.
@@
Given a node with k{{k}} out-neighbors, $\mathcal{T}$}} adds k{{k}}-2 extra nodes in $\mathcal{G}$$^{\prime}$ to create a right-branching binary tree. We denote by $r$ some root node and $v_{1},v_{2},\ldots v_{{k}}$ the out-neighbors of $v$ from $\mathcal{G}$}. $\mathcal{T}$}} introduces k{{k}}-2 extra nodes, $h_{1},h_{2},\ldots,h_{{{k}}-2}$, that are used as follows. The root node $r$ in $\mathcal{G}$$^{\prime}$ has edges to $v_{1}$ and $h_{1}$. For 0<i<k20<{{i}}<{{k}}-2, we instantiate the edges $(h_{{{n}}-2},v_{{{n}}-1})$ and $(h_{{{n}}-2},v_{{{n}}})$. For i=n2{{i}}={{n}}-2, we instantiate the edges $(h_{{{n}}-2},v_{{{n}}-1})$ and $(h_{{{n}}-2},v_{{{n}}})$. The resulting tree is a binary right-branching tree. Crucially, we get the following fact: If there is a path between two nodes $v_{1},v_{2}$ in $\mathcal{G}$}, there is a path between the corresponding nodes $v_{1},v_{2}$ in ${{\mathcal{T}}}({\mathcal{G}})$. Our construction trivially preserves reachability query results. We now show how to perform $\mathcal{T}$}} on transformers.